[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Jun 10, 2016

Dr. Craig A. Mattocks craig.mattocks at noaa.gov
Sun Jun 12 14:30:23 UTC 2016


The problem seems to have occurred at NESDIS. Once NCO switched from the NESDIS STAR ‘ftp2' server to the ‘ftp' server, the data began flowing again.

>From the SDM’s log:

HIMAWARI DATA FLAGS

3a. 1020Z - BB is showing purple ftp and red ingest/products
flags for the himawari-a1 system.  An email was sent to the STAR
folks about this.  IDP's Viktor was also called...he will
investigate switching to the backup STAR system. (MAS/GWN) 

3b. 1155Z - Viktor reported that he switched processing from the
ftp2 STAR system to the ftp STAR system and that we are now
processing Himawari data normally.  I see the same in BB and in
NMAP.  Viktor also reports it looks like the BB ftp test for the
a1/b1 systems is hung up on ftp2...that is why the flags went
purple...he will work on clearing these.

3c. 1311Z - Viktor was able to clear the ftp purple flag on the
primary b1 server.  However he cannot clear the purple flag on
the backup server at this time.  This monitoring issue on the
backup machine will be dealt with further on Monday.

Craig

> On Jun 12, 2016, at 10:00 AM, Mark DeMaria - NOAA Federal <mark.demaria at noaa.gov> wrote:
> 
> Chris:
> 
> The satellite wind processing on the compute server is maintained by TSB. The ssec server is not operationally supported by UW, but it usually is fairly reliable. Most likely it is an issue at Wisconsin, unless some firewall change prevented us from connected. Also, in the longer term, I would like to get the sat winds from an operational source at NESDIS. 
> 
> Mark
> 
> On Sat, Jun 11, 2016 at 11:16 AM, Christopher Mello - NOAA Federal <christopher.mello at noaa.gov> wrote:
> All,
>  I got an alarm indicating the sat winds are not updating...I t looks like a connection issue with  stratus.ssec.wisc.edu where the comp servers download the data...I will open up an NCO ticket...It is currently unclear what side of the firewall the problem originates from.
> 
> Chris 
> 
> 
> On Sat, Jun 11, 2016 at 12:42 AM, HSU Shift Log <nhc.tsbadmin at noaa.gov> wrote:
> HSU Shift Log for Sat Jun 11 2016 00:42:14 GMT-0400 (EDT)
> 
> Shift: R
> 
> Forecaster: M Brennan
> 
> Technical Issues Summary: 
> 
> ATCF issues: None
> 
> N-AWIPS/AWIPS2 issues: None
> 
> Other Techincal issues: None
> 
> 
> 
> -- 
> Chris Mello
> National Hurricane Center
> TSB
> 




More information about the Ncep.nhc.nco_contacts mailing list